OpenStudy (anonymous):

Solve 2x2 + x = 15. x = −5 and x = 5 x = three over two and x = −5 x = 15 and x = −2 x = five over two and x = −3

OpenStudy (anonymous):

-3 and 2.5

OpenStudy (anonymous):

Those aren't answers?

OpenStudy (anonymous):

Oh Nevermind

OpenStudy (campbell_st):

to solve this, rewrite it \[2x^2 + x - 15 = 0\] this can be factored or solved using the general quadratic formula for a quadratic \[ax^2 + bx + c = 0\] multiply a and c so you have 2 * -15 = -30 find the factors that add to 1, the larger factor is positive, so 6 and -5 next write it in factored form (ax + factor 1)(ax + factor 2) ------------------------- a so your problem becomes \[\frac{(2x + 6)(2x - 5)}{2} = 0\] remove any common factors \[\frac{2(x +3)(2x -5)}{2} = (x+3)(2x -5) = 0\] so you can solve x + 3 = 0 and 2x - 5 = 0 to get the solutions